you walk 4 miles at 30° and then 2 miles at 135°. what is the distance of your location from the starting point
Answer by stanbon(75887)

you walk 4 miles at 30° and then 2 miles at 135°. what is the distance of your location from the starting point
------
Draw the picture on a coordinate system.
Label the triangle ABC where A is the origin, AB is the 4 mile length,
BC is the 135 degree line, CA is the distance from origin to final position.
----
Sketch a vertical line thru B to the x-axis at point X.
----
Angle ABX has to be 60 degrees.
Angle ABC has to be 180-60-45 = 75 degrees.
---
Use the Law of Cosines to solve for CA.
---
CA^2 = 4^2 + 2^2 - 2*4*2cos(75) = 17.929
------------------
CA = 4.23 miles
-----------------------
Cheers,
Stan H.
=======================